OpenStudy (anonymous):

Simplify and show your work a^5 * a^4 --------- a²

jimthompson5910 (jim_thompson5910):

to simplify a^5 * a^4, you add the exponents

jimthompson5910 (jim_thompson5910):

a similar example would be x^3 * x^4 = x^(3+4) = x^7